Novocaine, used as a local anesthetic by dentists, is a weak base (Kb = 8.91 x 10^-6).

What is the ratio of the concentration of the base to that of its acid in the blood plasma (pH = 7.40) of a patient?


1
Expert's answer
2020-12-01T09:03:54-0500

The pH is known, and Kb is known, so calculate [-OH], then rearrange the Kb equation to get the ratio of the base to acid forms: [Novocaine]/[Novocaine•H+].

pH = 7.40, so pOH = (14 - 7.40) = 6.60 Therefore, [-OH] = 10-6.60 = 1.58 × 10-7 M

Rearrange:

Kb = [−OH][Novocaine⋅H+]/

[Novocaine]

to give:

     [Novocaine]/      

                                      =   [OH]      

[Novocaine⋅H+] Kb

So:

[Novocaine]/[Novocaine•H+] = 1.58 × 10-7 / 8.91 × 10–6

therefore, the ratio is = 1.8 × 10–2